# Scalable Private Search with Wally

### v2

@@ Abstract
This paper presents Wally  a private search system that supports efficient semantic and keyword search queries against large databases. When sufficient clients are making the queries, Wally performance is significantly better than previous systems. In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry. As a result, performance degraded proportionally with the number of entries in the database.
In Wally we get rid of this limitation. Specifically, for each query the server performs cryptographic operations only against a few database entries. We achieve these results by requiring each client to add a few fake queries, and sends each query via an anonymous network to the server at independently chosen random instants. Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake, Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases. Concretely, Wally can serve $8$M requests at a rate of 3,000 queries per second. That is around 60x higher than the state-of-the-art scheme.
@@ 1 Introduction
Consider a scenario where a client holding a search query wants to retrieve relevant information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query to respond accurately to each query, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers ”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
@@
Private search engines opt to provide query privacy. However, these systems have significantly higher computation and communication overhead than the state-of-the-art insecure solutions. For example, in Tiptoe [HDCZ23], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries. The high computation is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade off in a later section.. Additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
@@
In this paper, we present Wally, a privacy preserving search engine that efficiently scales to performance constraints of large-scale search engines. Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword. To provide privacy, our key observation is that in search engines, there are always enough clients making queries. We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries. In our protocol, the server computation and the communication get amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
@@
Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy. In Wally the server learns noised differentially private histogram of anonymized queries from all clients. We emphasize that Tiptoe, is fully oblivious, i.e., the server learns nothing besides that a particular client is making a query. Therefore, the privacy guarantee of Tiptoe is stronger than Wally  however differential privacy is an accepted standard for strong privacy [DR+14]. Specifically, in Wally every client is guaranteed that all the statistical inferences are insensitive to their data. In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
@@
At a high level, Wally works in epochs. Within each epoch, a fixed number of users make queries. The epoch size is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch. The server database is divided into $C$ clusters. Each query consists of a relevant cluster index, and the server response is all the entries in the cluster. The queries arrive at the server via an anonymous network, ensuring all identifiable information like IP address is removed from the queries. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network.
@@
To ensure that the server does not learn information by observing the access patterns, in Wally each client also makes fake queries. For each fake query, the client picks a randomly selected bucket. This ensures that the server cannot differentiate between real and fake queries. The number of fake queries is sampled from a particular distribution to ensure the revealed histogram of queries is $(\epsilon,\delta)$-differentially private. In expectation, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries where $U$ is the number of clients within an epoch, and $\Delta$ is the maximum number of queries each client can make in an epoch. Note that increasing the number of clients decreases the average number of fake queries. The overhead due to fake queries can be traded with privacy because more fake queries are required as $\epsilon$ decreases. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
Wally also ensures that arrival time and order of the queries leak no information about the client to the server. To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch. Also, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
@@
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in very high response overhead. In Wally  we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead. Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE. The server computes the distance function between the encrypted information and the cluster entries under SHE and returns encrypted scores. This reduces the response because the size of encrypted scores is significantly smaller than the entries. For a keyword search, the client includes encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword. To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data. We propose various optimizations that significantly improve performance for both kinds of searches.
@@
Wally requires decent number of clients always present in an epoch to make queries. We argue that this assumption aligns with current search engines, for example a hundred of millions of people make queries in Google search daily [Fle24]. We remark that for systems with few users, overhead due to fake queries could get quite high, therefore for these systems fully oblivious protocols like Tiptoe will be suitable.
@@ 2 Background
2.1 Homomorphic encryption
We use the BFV [Bra12, FV12] SHE scheme [^2] BFV is also a fully homomorphic scheme, which can homomorphically compute any function, with large enough parameter sets. We only focus on its simpler SHE parameter regime in our paper. where each plaintext is a polynomial in $\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and $t$ is a prime or prime power. Wally can also be implemented with the BGV [BGV14] SHE scheme with similar performance. Both schemes are post-quantum SHE schemes based on the ring-learning-with-errors (RLWE) problem [Reg05, LPR10]. A BFV ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})\in R_{Q}^{2}$ reduced modulo $Q$ and $X^{n}+1$ for a power of two $n$, $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$. The number $n$ is a ring dimension, $n\in\{2^{12},2^{13}\}$ in our use cases, and $Q$ is a ciphertext modulus with $\log_{2}Q=83$ or $\log_{2}Q=155$ since we use 128-bit secure post-quantum security parameters [ADPS16].
A BFV ciphertext satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=\lfloor Q/t\rceil m+e$ where $s$ is a secret key polynomial with small coefficients, in $\{0,\pm 1\}$, $e$ is small error polynomial, $t$ is the plaintext modulus, and $m\in\mathbb{Z}_{t}[X]/(X^{n}+1)$ is the plaintext polynomial. This error polynomial’s entries grow as we compute more homomorphic operations on the ciphertext and we can only decrypt if $\|e\|_{\infty}<Q/2t$. We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
@@ 2.1 Homomorphic encryption
Vectorized HE.
Vectorized plaintext spaces are used via polynomial interpolation: $X^{n}+1\mod t$ splits when $t$ is a prime satisfying $t=1\mod 2n$. In the vectorized case, each ciphertext encrypts a vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$ and the homomorphic operations are single-instruction multiple-data (SIMD) instructions. Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
@@
BFV’s $\mathsf{Eval}$ calls five subroutines:
@@
Ciphertext rotation and ciphertext multiplication require an evaluation key since the secret changes with each rotation and ciphertext multiplication. This operation is called “key-switching” for rotations and “relinearization” for ciphertext multiplications.
Table 1 has concrete run times and noise growth for BFV. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
@@
$\mathsf{CtCtAdd}$ $070$absent0$\approx$ 0≈ 0
$\mathsf{PtCtMult}$ $0.09$ $22$
$\mathsf{CtCtMult}$ $12.1$ $20$
$\mathsf{CtRotate}$ $3.6$ $\approx 0$
Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $8192$, the ciphertext modulus $Q$ has $150$ bits, and the plaintext modulus $t$ has $20$ bits. Time costs are measured with the SEAL library [CLP17] version 3.7.2 on a single core in a AWS t2.2xlarge instance.
Differential privacy
We use the standard notion of differential privacy (DP) [DMNS06, DKM+06]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private, if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
@@
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [Vad17, Lemma 2.3].
@@
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [GRS99], or mixnets [Cha81]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
Private information retreival.
Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation. Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database. PIR protocols can be for index retrieval or retrieval by keyword [CGN98]. In this paper, we focus on single-server PIR protocols [KO97] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard. We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [ACLS18] and [ALP+21]. There is a notion of differentially-private PIR in the multi-server setting [TDG16, AIVG22], where the database is stored across non-colluding servers while guaranteeing differential privacy. This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
3 System Design
Figure 1: High-level overview of Wally. Throughout an epoch, clients sends encrypted real and fake queries at random time slots to an anonymization network which “washes” queries of identifying information.
A high-level design of Wally is in Figure 1. It consists of three entities: the server, the clients, and the anonymization network. Wally runs in epochs of time; within each epoch, a fixed number of clients make queries. Wally does not require any synchronization across clients. Each client can decide to participate or skip an epoch. However, we require a minimum of clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. Overall, privacy loss is the sum of the loss due to each epoch in which the client has participated. The correctness guarantee is that all clients’ queries are processed by the end of the epoch. Wally’s privacy notion also requires that each client will make at most $M$ queries in the system’s lifetime; therefore, a can participate at max $M$ epochs.
Each epoch is divided into second-long time slots, and privacy is guaranteed by the following:

### v3

# Wally: An Efficient Private Search Engine

@@ Abstract
This paper presents Wally, a private search system that supports efficient semantic and keyword search queries against large databases. When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems. In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry. As a result, performance degraded proportionally with the number of entries in the database.
In Wally, we get rid of this limitation. Specifically, for each query the server performs cryptographic operations only against a few database entries. We achieve these results by requiring each client to add a few fake queries, and send each query via an anonymous network to the server at independently chosen random instants. Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake. Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases. Concretely, Wally can process eight million queries in just 39 mins. That is around four orders of magnitude less than the state of the art.
@@ 1 Introduction
Consider a scenario where a client holding a search query wants to retrieve relevant information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query in order to respond accurately, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
@@
Private search engines opt to provide query privacy. However, these systems have significantly higher computational and communication overhead than the state-of-the-art insecure solutions. For example, in Tiptoe [29], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries. The high computational cost is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade off in a later section.. An additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
@@
In this paper, we present Wally, a privacy preserving search engine that efficiently scales to the performance constraints of large-scale search engines. Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword. To provide privacy, our key observation is that in search engines, there are always enough clients making queries. We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries. In our protocol, the server computation and the communication are amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
@@
Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy. In Wally, the server learns a noised differentially private histogram of anonymized queries from all clients, over a partition of the server database. We emphasize that previous works, e.g., Tiptoe, are fully oblivious. That is, the server learns nothing besides that a particular client is making a query. Therefore, the privacy guarantee of schemes like Tiptoe is stronger than Wally. However, differential privacy is an accepted standard for strong privacy [19]. Specifically, every client in Wally is guaranteed that all the statistical inferences are insensitive to their data. In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
@@
At a high level, Wally works in epochs. Within each epoch, a fixed number of users make queries. The epoch length is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch. Server side, the database is divided into $K$ clusters. Each query consists of a relevant cluster index, and the server response is all the entries in the cluster. The queries arrive at the server via an anonymous network, ensuring all identifiable information like IP address is removed from the queries. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network.
@@
First, to ensure that the server does not learn information by observing the access patterns, each client in Wally makes fake queries. For each fake query, the client picks a randomly selected bucket. This ensures that the server cannot differentiate between real and fake queries. The number of fake queries is sampled from a particular distribution to ensure the revealed histogram of queries is $(\epsilon,\delta)$-differentially private. In expectation, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries where $U$ is the number of clients within an epoch, and $\Delta$ is the maximum number of queries each client can make in an epoch. Note that increasing the number of clients decreases the average number of fake queries. The overhead due to fake queries can be traded with privacy because more fake queries are required as $\epsilon$ decreases. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
Second, Wally also ensures that arrival time and order of the queries leak no information about the client to the server. To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch. Also, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
@@
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in very high response overhead. In Wally, we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead. Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE. The server computes the distance function between the encrypted information and the cluster entries under SHE and returns encrypted scores. This reduces the response because the size of encrypted scores is significantly smaller than the entries. For a keyword search, the client includes encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword. To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data. We propose various optimizations that significantly improve performance for both kinds of searches. Overall, a database with 35 million entries Wally can process eight million real queries in just 39 minutes on 320 cores set up, a drastic improvement over the state-of-the-art, which would have required years to process the same queries. Additionally, the size of each request is 53 times smaller and the size of each responses is 3.5 times smaller compared to Tiptoe.
@@
Due to space limitation, we discussed related works in Appendix A.
@@
Wally requires a set number of clients always present in an epoch to make queries. This assumption aligns with current search engines. For example, hundreds of millions of people make queries in Google search daily [22]. We remark that for systems with few users, the overhead due to fake queries could get quite high, therefore fully oblivious protocols, like Tiptoe, for these systems would be more suitable.
@@ 2 Background
2.1 (Somewhat) Homomorphic Encryption
Fully homomorphic encryption (FHE) is a special scheme allowing arbitrary computation over ciphertexts. FHE for arbitrary computation is still very expensive. To achieve practical performance, somewhat homomorphic encryption (SHE, also called leveled FHE) is often used, which only supports computation up to a fixed depth.
We focus on SHE schemes based on the Ring Learning with Errors (RLWE) problem. Concretely, Wally can be implemented using any SHE scheme such as BFV [21] or BGV [11] that share the following structure. A plaintext $m$ is a polynomial in a ring $R_{t}=\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and plaintext modulus $t$ is a prime or prime power. The secret key $s$ is a polynomial of degree $n-1$ with small coefficients, in $\{0,\pm 1\}$. A ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})=(a,as+m+e)\in R_{Q}^{2}$ where $Q$ is ciphertext modulus and $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$. Here $a$ is picked uniform randomly and $e$ is a noise polynomial with coefficients sampled from a bounded Gaussian distribution. A scheme satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=m+e$ This noise polynomial’s coefficients grow as we compute more homomorphic operations on the ciphertext, and we can only decrypt correctly if $\|e\|_{\infty}<Q/2t$. We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
@@
Vectorized SHE.
When a plaintext modulus $t$ is a prime satisfying $t=1\mod 2n$, these SHE schemes support operations on plaintext vectors as well as polynomials. In the vectorized case, each ciphertext encrypts a plaintext vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$, and the homomorphic operations are single-instruction multiple-data (SIMD) instructions, that is, the same operation performed by each component of plaintext vectors. Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
@@
SHE schemes we consider support following operations.
@@
In Table 1 we show concrete run times and noise growth for BFV, an efficient SHE scheme. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
Evaluation keys.
Ciphertext rotation and multiplication require evaluation keys that depend on a secret key. Additionally, each rotation requires a separate evaluation key. Usually, these keys can be reused, but As mentioned later, in Wally  the client must send fresh keys for each query. Therefore, these keys contribute to the client’s request size.
@@
$\mathsf{CtCtAdd}$ $0.004$ $0.5$
$\mathsf{PtCtMult}$ $0.02$ $20$
$\mathsf{CtCtMult}$ $2.5$ $26$
$\mathsf{CtRotate}$ $0.5$ $0.5$
Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $4096$, the ciphertext modulus $Q$ has $83$ bits, and the plaintext modulus $t=40961$ (16 bits). $\mathsf{CtRotate}$ noise is measured on repeated rotations and the others are from two input ciphertexts with the same noise levels. Time costs are measured with the swift-homomorphic-encryption library[7] (git commit b70d927) on an Intel Xeon w3-2423 using a single core.
Differential privacy.
We use the standard notion of differential privacy (DP) [18, 17]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private, if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
@@
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [40, Lemma 2.3].
@@
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [26], or mixnets [12]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
Private information retrieval.
Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation. Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database. PIR protocols can be for index retrieval or retrieval by keyword [14]. In this paper, we focus on single-server PIR protocols [33] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard. We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [5] and [4]. There is a notion of differentially-private PIR in the multi-server setting [38, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy. This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
3 System Overview
Figure 1: High-level overview of Wally. Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “washes” queries of identifying information.
A high-level design of Wally is in Figure 1. It consists of three entities: the server, the clients, and the anonymization network. Wally runs in epochs of time; within each epoch, a fixed number of clients make queries. Within each epoch, Wally performs a search that outputs the most relevant results to client queries, similar to insecure search engines. Wally does not require any synchronization across clients. Each client can independently decide to participate or skip an epoch. However, we require a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. Overall, privacy loss for a particular client is the sum of the loss due to each epoch in which the client has participated. Wally’s privacy notion also requires that each client will make at most $M$ queries in the system’s lifetime; therefore, a client can participate at max $M$ epochs. The correctness guarantee is that all the queries within epoch are processed by the end of the epoch.
In Wally Each epoch is divided into time slots, say a second long, and privacy is guaranteed by the following:
  • Hiding the content of the queries (somewhat homomorphic encryption).